On April 2, 2017, Victor, Inc. acquired a new piece of filtering equipment. The cost of the equipment was $480,000 with a residual value of $30,000 at the end of its estimated useful lifetime of 4 years.

a) Assume that in its financial statements, Victor uses straight-line depreciation and rounds depreciation for fractional years to the nearest whole month. Depreciation recognized on this equipment in 2017 and 2018 will be:

b) Assume that in its financial statements, Victor uses straight-line depreciation and the half-year convention. Depreciation recognized on this equipment in 2017 and 2018 will be:

c) If Victor uses straight-line depreciation with the half-year convention, the book value of the equipment at December 31, 2018 will be:
Answer #1

a) Depreciation expense per year = (480000-30000/4) = 112500

2017 Depreciation = 112500*9/12 = 84375

2018 Depreciation = 112500

b) Depreciation expense 2017 = 112500*6/12 = 56250

2018 Depreciation expense = 112500

c) Book value of equipment = 480000-112500-56250 = 311250